# Graph Linearization Methods for Reasoning on Graphs with Large Language Models

### v1

@@ 4 Dataset
Node Counting Max Degree Calculation
Motifs’ Shape
Classification
Accuracy MAE Accuracy MAE Accuracy
Linearization
Method
ZS OS ZS OS ZS OS ZS OS ZS OS
Default Node Labels CoreNumber 43.39 45.7 1.16 1.12 19.01 20.55 1.93 2.04 39.2 38
Degree 53.85 62.28 1.03 1.12 23.67 30.89 1.50 1.70 46.1 39.57
PageRank 50.92 51.38 2.09 1.9 20.34 29.92 1.67 1.63 47.27 33.5
Linegraph 41.59 34.71 1.52 1.92 19.21 11.74 1.77 2.05 44.63 33.5
Node Relabeling CoreNumber 42.1 28.39 0.9 1.55 13.70 17.51 2.09 1.95 41.43 46.9
Degree 57.70 54.58 0.66 0.74 24.64 28.32 1.45 1.38 44.13 43.03
PageRank 47.75 42.21 1.58 1.57 18.21 19.61 1.77 1.81 46.5 35.37
Linegraph 34.34 26.18 0.97 1.14 14.30 11.27 2.04 2.15 45 26.8
Baselines MotifAware 45.92 37.48 1.15 1.35 10.40 11.01 3.74 4.18 49.37 31.93
Random 48.95 41.42 0.6 0.8 11.55 8.80 2.55 2.60 43.95 31.89
Table 1: Comparison of various linearization methods across all tasks using Llama 3 Instruct 8B. The table presents the performance of different linearization methods, including the default node ID labeling against the node relabeling version, evaluated under both zero-shot (ZS) and one-shot (OS) settings. Metrics used for comparison are mean absolute error (MAE) and accuracy, where a lower MAE indicates more precise predictions, and higher accuracy reflects the model’s ability to correctly respond to the given task. Bold indicates the best performance in each setting, while underlined scores represent the second best method in those settings.
5 Experimental setup
In this section, we describe the protocol we followed to test the extent to which our linearization methods perform on graph reasoning tasks.
Tasks and prompts. Our experiments include three graph reasoning tasks. The first task is node counting, where the LLM infers the number of nodes from a linearized graph’s edge list. The second task is max degree calculation, where the LLM determines the maximum node degree. The third task, motifs’ shape classification, gives the LLM detailed definitions of the five motifs’ shape in our dataset, and ask it to identify which motif is present in a graph. This combination of reasoning and classification tasks allows us to comprehensively evaluate LLMs’ understanding of structural features and examine how different edge list orderings affect their performance.
@@ 5 Experimental setup
In every prompt, a node $i$ is represented by an incremental integer while an edge between nodes $i$ and $j$ is denoted by the bracketed pair $(i,j)$. An edge list is expressed as a sequence of edges, sorted according to the scheme used in each linearization method. We tested both zero-shot and one-shot approaches, where a randomly selected graph from the dataset was used consistently as the one-shot example across all experiments. The prompt templates are provided in Appendix A.
LLMs. For our experiments, we employed the 8B parameter version of Llama 3 Instruct (Dubey et al., 2024), quantized to 4 bits. We opted for $1\mathrm{e}{-3}$ temperature and $1\mathrm{e}{-1}$ sampling parameter values, in order to ensure more deterministic and coherent outputs, reducing randomness and enhancing consistency in generated responses. This allows us to assess its sensitivity to our linearization methods. The experiments were run on an NVIDIA A5000 GPU with 24GB of VRAM.
Baselines. For our comparisons, we used two baseline linearization methods.
MotifAware. This baseline uses the default ordering in which the graph is generated. As the graphs are built incrementally, starting with the base graph and then motifs are attached, this method produces a structure-aware edge list order (see Figure 1 “Default”). We expect this embedded structural knowledge to reflect on our tasks and increase LLM performance. However, generating such structure-aware lists requires knowledge of the graph construction process, which may not be feasible in real-world scenarios with larger, more complex graphs.
@@
Random. This baseline involves a fully Random ordering of the edge list, where edges are arranged without following any inherent scheme. To further eliminate structural biases, we also randomly shuffle the node IDs (labels). To mitigate the risk of skewed results, we applied five different random orderings and averaged their performance.
@@
Evaluation Metrics: We use exact accuracy to compare our methods, measuring the ratio of correct predictions as $\frac{1}{n}\sum_{i=1}^{n}I\left(y_{i}=\hat{y}_{i}\right)$, where $n$ is the number of graphs, $y_{i}$ the correct answer, and $\hat{y}_{i}$ the LLM’s response. For the counting tasks, we consider a result accurate only if it is an exact match. For the classification task, accuracy reflects the total across all shapes, requiring the predicted shape to appear at least once in the graph.
While accuracy provides a strict measure of correctness, it does not quantify the average magnitude of errors. To gain more fine-grained insights for counting tasks, especially when exact matches are not achieved, we also use Mean Absolute Error (MAE), calculated as $\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$, where $n$ is the number of graphs, $y_{i}$ the actual number of nodes (or max degree) and $\hat{y}_{i}$ the LLM’s response.
@@ 6 Experimental results analysis
In this section, we compare the performance across different tasks and linearization methods. The results are shown in Table 1.
Node counting. All methods show relatively low average errors. This suggests that the model is generally accurate in this task. However, models with lower average errors do not always show higher accuracy when compared. This indicates that while some methods frequently predict the correct node count, their errors are larger when they do occur.
Our proposed linearization methods consistently outperform the baselines in accuracy, with the Degree and PageRank methods showing the most potential. The Random baseline and the Relabeled Degree method exhibit nearly identical average errors, with one performing better in the one-shot setting and the other in the zero-shot setting. Although the Random baseline has a similarly low average error compared to the Relabeled Degree method, it suffers a significant drop in accuracy.
Notably, the node relabeling strategy leads to a performance drop in accuracy, except for the zero-shot Degree method. However, applying relabeling reduces the average error, except for the one-shot CoreNumber method. A consistent performance drop is observed in the one-shot setting across all methods. Finally, centrality-based methods tend to achieve higher accuracy than degeneracy-based approaches.
Max degree calculation. This task is more complex than the node counting task. Node counting only requires the model to identify the largest number in a sequence, whereas max degree calculation requires counting how often each node id appears and determining which id occurs most frequently. This additional layer of computation—counting and comparing occurrences—makes max degree calculation more challenging for LLMs. This complexity is reflected in the performance results across different setups.
With default node labeling, the Degree and PageRank methods stands out in one-shot setting, achieving the highest accuracy. The rest of the methods, while achieving competing average errors, maintain low accuracy scores.
When node relabeling is applied, the Degree method shows notable improvement in zero-shot, with increased accuracy and lower MAE. However, the other methods remain below 20% accuracy. Additionally, we observe that both baselines perform poorly, achieving very low accuracy.
These results highlight that Degree-based linearization is particularly effective for the max degree calculation task allowing it to handle the frequency-based complexity better than the others.
Motifs’ shape classification. For this task, we observe that the relabeling strategy has varied effects, leading to slightly higher accuracy scores across our methods mainly in the one-shot setting. However, similar to the node counting task, there is a consistent decline in performance from the zero-shot to the one-shot setting.
The MotifAware baseline stands out, performing exceptionally well in the zero-shot setting due to its embedded structural information, which likely enhances the model’s ability to recognize motif patterns. The PageRank method, however, performs on par. Unlike in the node counting task, centrality-based methods seem to yield accuracy comparable to that of the degeneracy-based ones.
@@ 7 Conclusion
Our research highlights the potential of large language models to extend their capabilities beyond text to encompass graph reasoning tasks. Central to this integration is the concept of “graph linearization”, where graphs are transformed into linear sequences of tokens, enabling LLMs to process them in a manner similar to natural language. We emphasize that effective graph linearization should preserve certain properties of natural language, such as local dependency and global alignment, to leverage the vast training LLMs have undergone with textual data. To this end, we developed and evaluated several graph linearization techniques grounded in graph centrality, degeneracy, and node relabeling strategies. Our experimental results on synthetic graphs underscore the superiority of these methods over random linearization approaches. This work paves the way for novel graph representations that are compatible with LLMs, fostering the integration of graph reasoning and machine learning within the broader scope of multi-modal processing in unified transformer models.
@@
Our study has a few limitations that future work could address. First, limited structural features were considered. We did not incorporate certain important graph features, such as communities and connected components, which may improve the method’s ability to capture complex structural patterns. Second, dataset diversity is limited. The experiments were conducted on a small set of synthetic graphs, so future research should explore a wider variety of graph datasets to ensure better generalizability. Third, task variety was narrow. While we focused on tasks like node counting and motif classification, future work could extend to more advanced tasks, such as graph clustering, matching, or even graph generation with fine-tuning.
Appendix A Prompt templates
Node Counting:

### v3

@@ 4.1 Datasets
GraphQA. Furthermore, we included GraphQA (Fatemi et al., 2023), which is widely used to evaluate the graph reasoning capabilities of LLMs. Similar to GraphWave, the dataset consists of randomly generated graphs derived from various graph generators, including Erdős–Rényi graphs (Erdős and Rényi, 1959), scale-free networks (SFN) (Barabási and Albert, 1999), the Barabási–Albert model (BA) (Albert and Barabási, 2002), and the stochastic block model (SBM) (Holland et al., 1983), along with star, path, and complete graph generators. A total of 500 graphs were sampled for ER, BA, SFN, and SBM models, whereas 100 graphs were sampled for path, complete, and star graphs due to their lower structural variability. All generated graphs contained between 5 and 20 nodes.
Node Counting Max Degree Node Degree Edge Existence Diameter Shortest Path Path Existence Motifs’ Shape Average
Random Labeling
CoreNumber 25.97 / 34.98 17.47 / 17.37 58.53 / 56.79 51.83 / 55.29 8.9 / 11.87 24.57 / 17.84 85.17 / 66.16 45.8 / 64.07 39.78 / 40.55
Degree 28.0 / 36.98 27.63 / 27.14 60.83 / 52.22 47.6 / 53.72 7.97 / 10.17 27.8 / 15.81 84.63 / 69.12 48.47 / 64.8 41.62 / 41.24
PageRank 28.81 / 39.18 24.10 / 23.57 59.4 / 56.42 44.9 / 47.52 8.83 / 11.27 27.37 / 16.34 84.63 / 72.69 44.93 / 65.73 40.37 / 41.59
LG{CoreNumber} 21.9 / 27.38 19.00 / 16.47 46.23 / 42.68 52.47 / 50.12 8.73 / 12.0 23.03 / 16.21 83.87 / 64.22 44.4 / 63.13 37.45 / 36.53
LG{Degree} 20.4 / 26.51 27.63 / 18.87 48.07 / 47.92 55.47 / 52.65 8.7 / 11.94 26.3 / 17.64 84.8 / 62.29 44.33 / 61.17 39.46 / 37.37
LG{PageRank} 29.1 / 27.28 18.83 / 17.97 46.8 / 40.61 47.9 / 49.02 8.47 / 12.14 24.73 / 16.04 87.07 / 70.52 42.4 / 60.0 38.16 / 36.7
Node Relabeling
CoreNumber 28.65 / 36.05 14.57 / 16.17 58.3 / 61.32 59.43 / 59.89 9.7 / 10.54 27.53 / 18.57 88.0 / 72.92 44.3 / 64.57 41.31 / 42.5
Degree 31.44 / 43.85 29.40 / 32.48 62.93 / 56.79 58.5 / 55.09 11.1 / 11.24 30.3 / 16.31 82.0 / 72.79 48.37 / 76.3 44.26 / 45.61
PageRank 34.35 / 38.01 26.07 / 25.24 65.37 / 56.92 50.83 / 46.08 10.47 / 11.17 32.33 / 17.67 82.13 / 72.89 47.6 / 68.63 43.64 / 42.08
LG{CoreNumber} 15.64 / 20.34 13.33 / 8.67 56.23 / 48.15 63.97 / 55.55 8.5 / 12.5 27.6 / 17.67 85.83 / 70.66 45.0 / 50.9 39.51 / 35.56
LG{Degree} 23.59 / 31.94 19.83 / 16.11 48.9 / 47.85 60.03 / 63.05 10.1 / 12.14 29.63 / 17.31 84.8 / 70.36 41.07 / 54.23 39.74 / 39.12
LG{PageRank} 27.57 / 37.45 17.03 / 18.64 51.8 / 48.45 54.37 / 51.15 9.73 / 12.34 26.6 / 15.41 85.47 / 71.72 39.4 / 51.77 39.0 / 38.37
Baseline 32.46 / 34.6 15.13 / 12.94 39.36 / 45.99 36.99 / 53.25 7.99 / 12.38 22.73 / 16.57 86.98 / 68.05 43.53 / 66.8 34.86 / 34.86
Table 1: Accuracy scores for all tasks on the GraphWave dataset using Llama 3 8B, including the overall average. Each task compares two node labeling schemes in zero-shot / one-shot prompting against the random Baseline. The linearization names represent the node ordering methods used (CoreNumber, Degree, or PageRank), while ‘LG{*} indicates graphs were transformed to the corresponding linegraph beforehand. Underlined scores denote the best-performing linearization method for each task, labeling, and X-shot combination; bold indicates task-wise best.
Node Counting Node Degree Diameter Motifs’ Shape
Random Labels
CoreNumber 76.47 / 75.29 72.47 / 63.75 3.97 / 4.07 59.1 / 61.43
Degree 78.2 / 80.96 73.93 / 63.69 2.43 / 4.57 58.07 / 61.4
PageRank 76.1 / 78.09 74.0 / 63.39 2.07 / 3.17 54.33 / 54.6
LG{CoreNumber} 77.8 / 77.36 68.83 / 57.02 4.77 / 6.97 46.8 / 56.23
LG{Degree} 78.9 / 81.16 72.23 / 62.65 4.57 / 7.5 51.5 / 54.2
LG{PageRank} 84.83 / 84.93 72.4 / 61.25 4.93 / 8.0 44.7 / 43.93
Node Relabeling
CoreNumber 89.1 / 86.06 75.87 / 70.99 8.47 / 11.3 54.97 / 53.4
Degree 84.83 / 82.53 76.0 / 69.39 7.27 / 11.8 58.23 / 56.57
PageRank 84.77 / 81.13 73.93 / 67.59 8.63 / 12.84 58.1 / 47.4
LG{CoreNumber} 76.27 / 76.93 75.47 / 65.32 8.77 / 12.84 50.9 / 52.37
LG{Degree} 87.23 / 84.46 76.63 / 68.66 8.47 / 14.67 48.83 / 46.3
LG{PageRank} 86.23 / 84.86 76.43 / 63.32 11.07 / 13.7 47.2 / 41.1
Baseline 83.24 / 83.04 69.69 / 59.81 4.52 / 7.72 41.57 / 42.5
Table 2: Accuracy scores for a subset of tasks on the GraphWave dataset using Llama 3 70B as an ablation study. Notations remain the same as in Table 1.
4.2 Tasks
Our experiments encompass a series of graph reasoning tasks, which can be broadly categorized into classification-based and numerical tasks. Numerical tasks, ranging in difficulty, require the model to produce a numerical output, either through structural computation or counting-based inference. This combination of tasks allows us to comprehensively evaluate LLMs’ understanding of structural features and examine how different edge list orderings affect their performance.
A fundamental task is Node Counting, where the LLM estimates the number of nodes. In Node Degree calculation, the LLM determines the degree of a node. A more advanced variant, Maximum Degree calculation, requires the LLM to internally calculate the degree of all nodes and then identify the maximum among them.
Beyond node-related tasks, we assess the ability to infer relational properties. In Edge Existence and Path Existence tasks the LLM is given a randomly selected pair of nodes and must determine whether an edge or a connecting path exists between them, respectively. In the Shortest Path task, the model must compute the length of the shortest path between two given nodes, requiring a deeper understanding of graph connectivity. The Diameter Estimation task requires the model to determine the longest shortest path in the graph, showcasing global graph structure understanding.
Finally, we evaluated Motifs’ Shape classification, a dataset-specific task leveraging GraphWave’s embedded structures, where the LLM is given definitions of the five motif types and asked is to identify which is present.
@@ 4.2 Tasks
In every prompt, a node $v$ is represented by an incremental integer, while an edge between nodes $v$ and $u$ is denoted by the bracketed pair $(v,u)$. An edge list is expressed as a sequence of edges, sorted according to the scheme used in each linearization method. We tested both zero-shot and one-shot approaches, where a randomly selected graph from the dataset was used consistently as the one-shot example across all experiments. The prompt templates are provided in Appendix D.
4.3 LLMs
We used the 8B parameter Llama 3 Instruct (Dubey et al., 2024) with a temperature of $1\mathrm{e}{-3}$ and a sampling parameter of $1\mathrm{e}{-1}$ for more deterministic outputs to assess sensitivity to our linearization methods. Experiments were conducted on an NVIDIA A5000. Additionally, we conducted partial experiments with the 70B-parameter model to evaluate its impact on graph reasoning capabilities and to verify the consistency of our approach. Finally, to explore the impact of model family differences, we also include experiments over Qwen 2.5 14B-1M Yang et al. (2025) in the Appendix A.
4.4 Baselines
@@
For our comparisons, we consider only a random baseline. This baseline involves a fully random ordering of the edge list, where edges are arranged without following any inherent scheme. To further eliminate structural biases, we also randomly shuffle the node labels. This baseline is founded on the fact that we are working with general graphs, where default labels or ordering are neither predetermined nor necessarily provide meaningful information in real-world applications. In addition, to mitigate the risk of skewed results, we applied five different random orderings and averaged their performance. Our random ordering can be considered comparable to prior studies, which tend to preserve the inherent structure of the generator.
Node Counting Max Degree Node Degree Edge Existence Diameter Shortest Path Path Existence Average
Random Labels
CoreNumber 60.77 / 24.28 15.88 / 18.67 54.44 / 37.23 70.68 / 61.59 3.28 / 18.53 49.21 / 57.25 95.51 / 98.39 49.97 / 45.13
Degree 62.18 / 28.91 25.6 / 29.75 54.99 / 29.4 68.02 / 58.6 3.39 / 14.82 50.24 / 53.34 94.39 / 98.71 51.26 / 44.79
PageRank 62.12 / 28.31 26.03 / 28.94 56.23 / 28.05 68.39 / 59.26 3.16 / 15.25 29.34 / 47.3 94.51 / 98.68 48.54 / 43.68
LG{CoreNumber} 61.14 / 23.94 10.81 / 25.89 49.09 / 36.65 69.26 / 58.83 2.9 / 16.97 42.08 / 42.35 95.92 / 98.71 47.31 / 43.33
LG{Degree} 59.53 / 27.22 12.97 / 28.25 48.46 / 37.69 68.25 / 60.39 2.7 / 16.34 23.99 / 43.1 95.2 / 98.82 44.44 / 44.54
LG{PageRank} 60.66 / 26.24 11.59 / 28.65 47.51 / 38.43 69.49 / 62.66 2.67 / 17.43 40.78 / 45.6 94.31 / 98.79 46.72 / 45.4
Node Relabeling
CoreNumber 61.06 / 34.38 22.09 / 30.29 51.94 / 30.49 72.02 / 70.17 2.36 / 14.41 46.28 / 52.85 94.79 / 98.56 50.08 / 47.31
Degree 65.31 / 45.48 39.11 / 44.3 54.76 / 29.2 70.18 / 74.6 2.93 / 14.3 46.22 / 52.01 96.89 / 98.76 53.63 / 51.24
PageRank 64.71 / 44.97 40.72 / 43.15 58.77 / 28.65 69.97 / 70.83 3.57 / 14.13 26.17 / 49.54 97.04 / 98.82 51.56 / 50.01
LG{CoreNumber} 63.56 / 32.54 22.69 / 33.49 47.48 / 29.14 72.91 / 59.15 1.93 / 15.1 24.33 / 43.84 96.52 / 98.82 47.06 / 44.58
LG{Degree} 64.77 / 35.07 27.87 / 35.27 47.2 / 39.18 70.46 / 59.26 3.19 / 15.13 42.45 / 46.72 96.06 / 98.76 50.29 / 47.06
LG{PageRank} 69.37 / 32.42 26.37 / 36.02 49.5 / 38.9 70.84 / 63.98 3.13 / 15.19 41.53 / 47.44 95.48 / 97.96 50.89 / 47.42
Baseline 66.28 / 27.38 9.28 / 20.37 49.78 / 35.86 67.02 / 58.99 2.94 / 14.98 37.7 / 44.55 95.83 / 98.51 46.98 / 42.95
Table 3: Accuracy scores for all tasks on the GraphQA dataset using Llama 3 8B, , including the overall average. Notations remain the same as in Table 1.
4.5 Evaluation
@@
We use exact accuracy to compare our methods, measuring the ratio of correct predictions as $\frac{1}{n}\sum_{i=1}^{n}I\left(y_{i}=\hat{y}_{i}\right)$, where $n$ is the number of graphs, $y_{i}$ the correct answer, and $\hat{y}_{i}$ the LLM’s response. For numerical tasks, we consider a result accurate only if it is an exact match. For the motifs’ shape classification task, accuracy reflects the total across all shapes, requiring the predicted shape to appear at least once in the graph.
@@ 5 Experimental Results Analysis
In this section, we review the performance of our linearization methods across various tasks.
Tables 1 and 3 present the performance of our methods on the GraphWave and GraphQA datasets with the Llama 3 8B model. Additionally, to evaluate model scale, we report results for selected GraphWave tasks using the 70B model in Table 2. The results are organized into three groups: one where node labels in the linearized graphs are randomly assigned, ensuring a fair comparison since, in real-world graphs, labels might be arbitrary; another where node labels are reindexed according to each method, as described in Section 3; and finally, a comparison against the random baseline. The performance related to the pseudo-random (default) node labels originally provided by synthetic graph generators is discussed in Appendix A.
Overall, across both datasets, our linearization methods consistently outperform the random baseline, highlighting the critical role of graph linearization, as evident in the average performance and across multiple individual tasks. Notably, on the GraphQA dataset, the combination of degree-based ordering and node relabeling improves performance by approximately 35% on the maximum node degree estimation task and by around 13% on the shortest path task. Similarly, on the GraphWave dataset, the combination of degree-based ordering and node relabeling enhances edge existence performance by roughly 26%. Comparing random and node relabeling reveals that ordering alone offers significant improvements over the baseline, while the additional information from structured relabeling further enhances accuracy on nearly all tasks. Among all tasks, diameter estimation is the most challenging across both datasets, with consistently low performance, indicating LLMs struggle to infer global graph properties at this level of complexity.
Linegraph-based methods (LG{*})—where edges are reinterpreted as nodes—highlight the importance of edge-to-edge relationships. While their overall average score is lower, they generally perform better in edge-based tasks, such as edge existence and path reasoning, by capturing interdependencies that might be less evident in traditional node-focused representations. These findings suggest that a more suitable linearization approach may be necessary to fully exploit the benefits of the linegraph transformation.
Similarly, CoreNumber-based methods achieve better performance in edge-centric tasks, which can be attributed to its ability to capture the structural cohesiveness of a graph. By emphasizing nodes embedded in densely connected subgraphs, core number ordering effectively preserves key connectivity patterns, making it particularly advantageous for reasoning about edge relationships. In contrast, while Degree- and PageRank-based orderings demonstrate the most consistent performance across various tasks, their strengths are more pronounced in node-related tasks.
When moving from zero-shot to one-shot setting, we notice a performance loss on binary classification tasks like edge and path existence. This decline may result from the model’s reliance on a single graph example, which does not fully capture the complexity and diversity of the dataset. However, despite this drop, one-shot prompting remains effective for more complex tasks.
Finally, when comparing the results between the 8B and 70B versions of Llama, we observe a significant performance boost in the node counting and motif shape classification tasks with the larger model. In contrast, the node degree task shows only a modest improvement. Meanwhile, diameter estimation continues to exhibit very low accuracy, indicating that this task remains particularly challenging regardless of model size.
@@ 6 Conclusion
In this study, we investigated different graph linearization techniques for LLMs. The core challenge lies in converting graphs into linear token sequences while maintaining important structural features—like local dependencies and global coherence—akin to those in natural language, to help LLMs more effectively interpret graph-based data. To this end, we developed and evaluated several linearization methods based on graph centrality and degeneracy. Our experiments showed that graph linearization notably enhances the performance of LLMs on graph reasoning tasks, especially when paired with the node relabeling technique. Our work presents novel graph representations tailored for LLMS, paving the way for integrating graph machine learning with the growing trend of multimodal processing through a unified transformer.
@@
Our study has several limitations that future research could address. First, we considered only a limited set of structural features. Key graph properties, such as community structures and connected components, were not incorporated, yet they could enhance the model’s ability to capture complex structural patterns. Second, the diversity of datasets is limited. We conducted our experiments exclusively on synthetic graphs, and future work should explore a broader range of graph datasets to improve real-world generalization. Third, the range of LLMs evaluated was narrow. We primarily focused on LLaMA 3, and future studies could investigate a wider variety of models, including those fine-tuned for coding or mathematical reasoning, to assess their impact on graph-based tasks. Fourth, our method inherits a fundamental limitation from the underlying LLMs: the input sequence is bounded by the model’s context window. This imposes an upper limit on the number of edges that can be represented per graph. In Appendix C, we analyze how this constraint translates to a maximum graph size under different edge densities and token budgets. Lastly, our study does not investigate the impact of additional training to adapt models to our linearization methods, which could potentially enhance their graph reasoning capabilities.
Appendix A Additional Experiments
Pseudo-random linearization. We also investigated the performance of utilizing the edge ordering directly provided by the graph generator. Most non-random graph generators create graphs procedurally, inherently embedding structural information within the edge list order. For instance, in GraphWave, the process begins with a base graph and subsequently attaches motifs, making it easier to distinguish between different structures. We hypothesize that this embedded structural knowledge will enhance task performance and boost the LLM’s capabilities. However, generating such structure-aware edge lists requires an understanding of the graph construction process, which may not be feasible for real-world applications involving larger and more complex graphs.
The results of both datasets are presented in Table 4. For the GraphWave dataset, the default edge ordering shows mixed performance compared to the random ordering baseline. When combined with structured edge ordering (Table 1), accuracy improves consistently across tasks, except for path existence. When comparing default labeling with structured node labeling, performance generally improves, though default labeling remains competitive in certain tasks. For the GraphQA dataset, the default edge ordering performs significantly better than the random ordering across all tasks except path existence. In this case, the default ordering proves to be particularly robust, making it challenging for structured edge ordering to achieve higher accuracy. Even compared to structured edge ordering (Table 3), default ordering often maintains a performance advantage, highlighting its effectiveness in this dataset.
Node Counting Max Degree Node Degree Edge Existence Diameter Shortest Path Path Existence Motifs’ Shape
Default Labels
CoreNumber 25.06 / 37.41 17.87 / 16.67 60.9 / 62.45 54.43 / 57.69 7.93 / 11.74 29.97 / 20.74 83.1 / 70.32 43.43 / 63.23
Degree 32.66 / 51.18 28.37 / 26.24 60.77 / 57.79 55.5 / 49.22 8.97 / 11.54 31.7 / 15.97 78.9 / 67.36 47.8 / 61.4
PageRank 35.77 / 49.38 24.40 / 26.88 63.43 / 56.42 49.4 / 53.08 8.63 / 11.5 31.33 / 18.24 81.3 / 73.99 48.03 / 65.27
LG{CoreNumber} 19.79 / 22.21 21.60 / 13.40 55.37 / 47.75 59.4 / 53.35 10.07 / 11.9 27.03 / 19.74 79.4 / 73.06 44.43 / 56.23
LG{Degree} 24.6 / 32.21 23.23 / 17.84 49.33 / 41.35 56.4 / 52.72 10.67 / 11.04 29.33 / 19.47 79.67 / 71.76 43.87 / 52.73
LG{PageRank} 34.83 / 41.78 17.90 / 19.14 54.3 / 50.18 48.9 / 52.15 8.77 / 12.17 27.63 / 16.47 79.77 / 76.59 43.07 / 48.63
Default Ordering 36.75 / 44.38 9.77 / 10.17 47.83 / 41.18 37.23 / 47.48 7.3 / 11.5 30.3 / 16.67 85.5 / 82.79 54.63 / 55.73
Node Counting Max Degree Node Degree Edge Existence Diameter Shortest Path Path Existence
Default Labels
CoreNumber 61.23 / 38.52 22.4 / 34.35 52.72 / 44.51 72.19 / 68.07 1.58 / 9.29 46.65 / 58.57 94.62 / 97.55
Degree 69.31 / 56.88 44.32 / 52.19 54.53 / 26.64 70.23 / 70.4 2.36 / 12.31 25.48 / 54.2 97.12 / 98.79
PageRank 69.03 / 53.97 44.23 / 52.1 54.56 / 33.8 69.66 / 71.66 2.88 / 11.94 45.41 / 50.86 96.98 / 98.79
LG{CoreNumber} 67.73 / 38.38 22.66 / 30.09 47.63 / 33.75 71.87 / 61.71 1.27 / 11.68 40.44 / 47.07 96.69 / 98.79
LG{Degree} 66.9 / 42.84 26.43 / 34.75 46.85 / 38.67 71.99 / 62.49 2.3 / 14.38 41.39 / 46.38 95.8 / 98.82
LG{PageRank} 70.87 / 37.46 24.82 / 35.13 49.5 / 37.83 72.36 / 58.92 2.24 / 16.4 41.53 / 47.61 95.17 / 97.93
Default Ordering 68.43 / 51.06 52.78 / 48.48 56.08 / 33.08 67.33 / 71.06 4.43 / 30.72 28.5 / 41.94 98.25 / 95.51
Table 4: Accuracy scores for all tasks on the GraphWave (top) and GraphQA (bottom) datasets using Llama 3 8B. For each task, we compare the default labeling scheme, as provided by the corresponding graph generator, against the default order the edges have been generated. Notations remain the same as in Table 1.
LLM Family Variation with Extended Context. To further examine the influence of model architecture and extended context capacity, we evaluated the performance of Qwen 2.5 14B-1M  Yang et al. (2025), a large-context language model from a distinct model family capable of processing input sequences of up to 1 million tokens. This evaluation allows us to assess whether architectural differences impact performance on graph-based reasoning tasks. Table 5 reports the results obtained for the same selection of tasks previously used in Table 2, enabling a direct comparison across models with varying capacities.
Node Counting Node Degree Diameter Motifs’ Shape
Random Labels
CoreNumber 70.67 / 72.12 76.5 / 68.66 12.1 / 12.5 58.5 / 56.3
Degree 71.23 / 70.96 76.97 / 66.66 12.07 / 13.3 62.27 / 59.43
PageRank 69.77 / 66.99 79.53 / 72.09 12.77 / 12.04 56.9 / 58.5
LG{CoreNumber} 71.13 / 68.92 72.17 / 54.08 12.63 / 14.17 59.17 / 61.97
LG{Degree} 71.77 / 72.86 74.8 / 62.49 13.97 / 15.37 60.5 / 60.6
LG{PageRank} 76.53 / 76.59 72.07 / 65.19 13.07 / 12.44 51.9 / 52.77
Node Relabeling
CoreNumber 86.73 / 87.66 80.23 / 75.49 12.93 / 14.5 58.03 / 53.17
Degree 74.2 / 77.63 78.03 / 72.59 12.8 / 11.3 61.3 / 49.83
PageRank 73.4 / 75.23 79.3 / 73.66 14.33 / 13.14 59.73 / 49.83
LG{CoreNumber} 76.03 / 71.52 78.77 / 69.19 12.63 / 12.64 64.03 / 57.53
LG{Degree} 80.03 / 81.39 76.6 / 69.22 15.4 / 12.77 60.57 / 46.77
LG{PageRank} 84.0 / 84.29 74.8 / 67.96 14.47 / 12.34 52.53 / 44.2
Baseline 78.84 / 77.43 70.09 / 69.19 9.95 / 10.74 54.53 / 57.3
Table 5: Accuracy scores for a subset of tasks on the GraphWave dataset using Qwen 2.5 14B - 1M as an ablation study. Notations remain the same as in Table 1.
Although scaling from LLaMA 8B to 70B yields substantial gains in tasks such as node counting and motif shape classification, Table 5 demonstrates that Qwen 2.5 14B—despite having fewer parameters than LLaMA 70B—achieves competitive, and in some cases superior, performance across several tasks. This is particularly evident in motif shape classification and diameter estimation, where Qwen’s results rival or exceed those of the larger model. Nonetheless, in line with trends observed across LLaMA variants, diameter estimation remains a consistently challenging task, with overall accuracy remaining low regardless of model architecture or scale.
Appendix B Performance and Complexity
All considered graph measures, such as core number, degree, and PageRank, are computationally efficient. For graphs where the number of edges exceeds the number of nodes, the computational complexity scales linearly with the number of edges, that is, $\mathcal{O}(m)$, where $m$ denotes the number of edges. Given that the graphs in the evaluated datasets are relatively small, the computation time required for these measures is negligible compared to the response generation time of the LLMs.
To illustrate this, we report on Table 6 the number of tokens generated per task along with the average time taken to produce a single response using LLama 3 8b. Numerical tasks, LLM responses are concise and rapidly converge to a final answer. In contrast, more complex tasks elicit longer responses that often involve intermediate reasoning steps.
Task Number of Tokens Inference Time (sec)
Node Counting 16 0.6
Max Degree 16 0.6
Node Degree 16 0.6
Edge Existence 128 4.8
Diameter 128 4.8
Shortest Path 128 4.8
Path Existence 128 4.8
Motif’s Shape 16 0.6
Table 6: Tokens generated and response time for a single graph using LLaMA 3 8B across tasks.
Appendix C Graph Size Limitations
In our approach, the entire graph is linearized into a token sequence and embedded directly into the model’s input prompt. As a result, the maximum size of the graph that can be processed in a single prompt is constrained by the model’s context window. Since the sequence length is primarily determined by the number of edges, we estimate the maximum number of edges that can be encoded per prompt for each model considered in this study.
To compute these estimates, we assume that each edge requires approximately 5 tokens to represent, and that the accompanying task description consumes an average of 100 tokens. Under these assumptions, Table 7 reports the estimated edge capacity corresponding to the context window of each model.
LLM Context Length (tokens) Max Number of Edges
LLama 3 8b 8,192 1,618
LLama 3 70b 8,192 1,618
Qwen 2.5 14B 1M 1,010,000 199,980
Table 7: Estimated Maximum Number of Edges considering Context Window Size.
In Table 8, we present statistics for several widely used graph datasets. Many of these graphs are sufficiently small to fit within the context window of contemporary LLMs, with notable exceptions such as large-scale social and e-commerce networks (e.g., Reddit, Amazon). This indicates that a substantial portion of benchmark graph datasets can be fully serialized and input to an LLM in a single prompt. Nevertheless, in practical applications, even graph neural networks often rely on sampling strategies rather than processing entire graphs at once. A similar strategy may be necessary when using LLMs for real-world graph tasks, depending on the application and scale.
Although our study primarily investigates the ability of LLMs to understand and reason over complete graph structures, we recognize that some of the tasks examined—such as node counting—are primarily diagnostic and may have limited practical relevance. These tasks are intended to serve as controlled benchmarks to assess the reasoning capabilities of LLMs, rather than to reflect typical graph processing workloads.
Dataset (Source) #Nodes #Edges
Cora Kipf and Welling (2016) 2,708 1,433
Citeseer Kipf and Welling (2016) 3,327 9,104
PPI Zitnik and Leskovec (2017) 2,245 61,318
PubMed Kipf and Welling (2016) 19,717 88,648
Amazon Yang et al. (2018) 12,752 491,722
OGBN-Proteins Hu et al. (2020) 132,534 39,561,252
Reddit Hamilton et al. (2017) 232,965 114,615,892
Amazon Products Chiang et al. (2019) 1,569,960 264,339,468
Table 8: Node and edge counts for commonly used graph datasets.
Appendix D Prompt templates
Node Counting